OpenStudy (anonymous):

Differentiate y= ln l 1+t-t^3 l

OpenStudy (helder_edwin):

is there an absolute value inside the natural logarithm?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

cool

OpenStudy (anonymous):

what do u do?

OpenStudy (anonymous):

just 1/ (1+t-t^3)?

OpenStudy (helder_edwin):

no. first u have to find where the polynomial is positive. in the points where the polynomial changes sign the derivative won't exist, because there we shal have corners.

OpenStudy (anonymous):

so..i need to sketch a graph?

OpenStudy (helder_edwin):

i would first solve \[ \large 1+t-t^3=0 \]

OpenStudy (helder_edwin):

the sketch of the graph would help, of course.

OpenStudy (anonymous):

solve u mean find derivative?

OpenStudy (anonymous):

but my book has a rules that d/dx lnlxl = 1/x

OpenStudy (helder_edwin):

then all u have missing is the chain rule

OpenStudy (helder_edwin):

by the way. what book is it?

OpenStudy (helder_edwin):

just a remark about what your book says: if we truly had \[ \large \frac{d}{dx}\ln|x|=\frac{1}{x} \] then the slope of the tangent line at \((|-1|,0)=(1,0)\) would be \(m=-1\) so the tangent line would be \[ \large y=-(x-1) \]

OpenStudy (helder_edwin):

http://www.wolframalpha.com/input/?i=graph+ {y%3Dln+x%2Cy%3D-%28x-1%29}

OpenStudy (helder_edwin):

what do u think?